OpenStudy (anonymous):

Two buses leave towns 616mi apart at the same time and travel toward each other. One bus travels 8mi/h slower than the other. If they meet in 4 hours, what is the rate of each bus?

OpenStudy (anonymous):

let the speed of the busses be x and x+8 mi/h the n since they move in opposite direction relative speed =x+x+8=2x+8 now by the problem (2x+8)*4=616 or 2(x+4)*4=616 or (x+4)=616/8 x+4=77 x=73 thus their speed are 73 and 81